tax:

Another Algebra 2 question: 2nd one out of 5

tax:

\[x^4 - 5x^2 - 36 = 0\] Here's the problem My final answer was: \[x = 3, -3, 2i, -2i\] Or in other words: 4 solutions

xXAikoXx:

It seems your answer is accurate if you ask me. Here's my solution: \[u^2 - 5u - 36 = 0\] \[(u-9)(u + 4) = 0\] u = 9... you can also say... u = -4 Finally we substitute \ \[x^2=9⇒x=±3\] \[x^2=−4⇒x=±2i\] Yes, you are 100% correct on this

xXAikoXx:

there's an equation missing because it didn't let me type in the 'equation' thing but its fine because it's not really relevant to this problem but yeah (:

tax:

@xxaikoxx wrote:
there's an equation missing because it didn't let me type in the 'equation' thing but its fine because it's not really relevant to this problem but yeah (:
Okay! Thank youu

yhyhR:

why you guys so smart bro!